OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3, how do I find the other two coordinates!?

OpenStudy (across):

What have you tried?

OpenStudy (zzr0ck3r):

Our equation will be in the form \(y=mx+b\) where \(b\) is the y value of the y-intercept and \(m\) is slope. they gave you slope so we have \(y=3x+b\), so we need to find \(b\) So we use the point \((4,2)\). Plug this in to the equation \(2=3(4)+b\) Solve for \(b\).

OpenStudy (anonymous):

Would it be 6? Probably not. What exactly do I do to solve for b in this "2 = 3 (4) + b"?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

\(2=12+b\) subtract \(12\) from both sides.

OpenStudy (anonymous):

y = 3x - 10?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Thank you, you're awesome!

hero (hero):

Another way to do it is to use the slope formula: \[\frac{2 - y}{4 - x} = 3\] \[2 - y = 3(4 - x)\] \[2 - y = 12 - 3x\] \[-y = 10 - 3x\] \[y = 3x - 10\]

OpenStudy (anonymous):

Why thank you for showing me this also, kind sir. I really appreciate the help I have received.

hero (hero):

You're welcome